OpenStudy (anonymous):

What is the diameter of the following circle: (x + 4)2 + (y – 9)2 = 18

OpenStudy (anonymous):

diameter =2r

OpenStudy (mathstudent55):

Circle \((x - h)^2 + (y - k)^2 = r^2\) has radius r and diameter 2r.

OpenStudy (anonymous):

so times 18 by 2?

OpenStudy (mathstudent55):

No. 18 is r^2. First find r. Then multiply it by 2.

OpenStudy (anonymous):

im confused

OpenStudy (freckles):

is is another word for equal r^2=18 <--can you solve this for r so that you have the radius

OpenStudy (anonymous):

so r is 9 ?

OpenStudy (freckles):

hmm 9^2=9(9) isn't 18

OpenStudy (mathstudent55):

I'll give you an example: In the circle \((x - 4)^2 + (y - 3)^2 = 9\), the center is (4, 3), and the radius is 3, since 3^2 = 9. You take the square root of 9 and get 3. That is the radius.

OpenStudy (mathstudent55):

In your case. r^2 = 18 Take the square root of 18 to find the radius. Then multiply the radius by 2 to find the diameter.

OpenStudy (anonymous):

so it'll be 8.48?

OpenStudy (mathstudent55):

You probably need an exact answer. This is how you do it. \(r = \sqrt{18}\) \(d = 2\sqrt{18} = 2 \sqrt{9 \cdot 2} = 2 \sqrt{9}\sqrt{2} = 2 \cdot 3\sqrt{2} = 6\sqrt2\)